The bottom of a 13 foot straight ladder is set into the ground 5 feet away from a wall. When the top of the ladder is leaned against the wall, what is the distance above the ground it will reach?

Answer by checkley77(12844)

A^2+B^2=C^2
5^2+B^2=13^2
25+B^2=169
B^2=169-25
B^2=144
B=SQRT144
B=12 FEET ANS.